OpenStudy (anonymous):

HELP! simplify: cos(180° + x) + cos(180° - x)

OpenStudy (sidsiddhartha):

cos(180° + x)=-cosx cos(180° - x)=-cosx so cos(180° + x) + cos(180° - x)=-2cosx

OpenStudy (anonymous):

perfect can you also help with this one? cos(126°)cos(36°) + sin(126°)sin(36°)

OpenStudy (sidsiddhartha):

use cos(A-B)=cosAcosB+sinAsinB

OpenStudy (anonymous):

i know im just not sure how!

OpenStudy (sidsiddhartha):

just using the formula-- cos(126°)cos(36°) + sin(126°)sin(36°)=cos(126-36)=cos90=0

OpenStudy (anonymous):

thank yoU!

OpenStudy (sidsiddhartha):

yw!!